Green Bank, WV
March 20,  2002

TO: J. Acree, C. Beaudet, R. Fisher, J. Ford, F. Ghigo, R. Lacasse,  R. McCullough, C. Niday, W. Sizemore
FROM: R. Fisher
SUBJ.: Minutes of Interference Protection Group 3-19-02 Meeting

In attendance: Acree, Beaudet, Fisher, Ford, McCullough, Sizemore

Minutes of Previous Meeting

 Status of current action items:

        1 - Reported-RF tracking data base (Jeff)

Jeff's data base remains as described at the last IPG meeting. John is working on an SQL version that will be accessible from the web. This is really two linked data bases. One is a list of all known RFI, and the other contains dossiers of signal characteristics.
        2 - Inventory of on-site antennas (Wes)
An email survey turned up six or eight antennas around the site that need to be checked for unintentionally radiated signals coming from the receivers to which they are connected. Carla mentioned that there are a number of wire antennas connected to boom-boxes that should be investigated. She will supply a list.
        3 - 300-400 MHz search and clean-up (Jeff, Carla)
Many sources of RFI were documented by John and Carla in their campaign to measure and turn off computers and other digital equipment in the lab one evening. However, turning off as much equipment as possible in the lab did not make much of a dent in the RFI spectrum radiating from the lab as seen on the deer stand antenna. Turning off equipment in the OVLBI trailer did eliminate a number of signals seen at the deer stand. It could be that the lab shielded rooms have sprung a leak or that the equipment in these rooms radiates more than can be suppressed by the rooms' 60 dB attenuation. Roger and Chuck have been asked to measure the shield's current effectiveness. Carla is working with Mike Holstine and the phone company on adding a good earth ground to the PBX system by drilling through the basement concrete where the PBX is housed. The unshielded LAN system was also found to be a significant source of RFI in the lab survey. John is working with Ed Childers on this.
        4 - Interference to pulsar observing around 1400 MHz
No progress.
        5 - Feed arm servo (Rich, Randy)
Randy reported that the shield box and plate are out of the shop, and that he will measure its effectiveness in the coming weeks. A design for the motor filters is essentially complete with as much as 120 dB of conducted signal attenuation expected down to 40 MHz. Thirty six filters at $250 each are required. Six of these have been put on a specially expedited order ($250 surcharge) to get them for initial tests. The first six should arrive in four weeks.
        6 - Lasers (Carla, Randy)
Two Hoffman boxes have been acquired for shielding purposes. The big one will house the main control circuitry, and the little one is intended to shield the 1500 MHz signal from each laser unit. The laser units will require quite a bit of work because there are a number of unfiltered wires in the current unit that are acting as 1500 MHz antennas. Negotiating a fix acceptable to the metrology group is taking time.
        7 - Boltec lightning detector (Chuck, Carla)
Carla reported that Chuck has determined what signals are being radiated by the lightning receiver unit and he is working on shields and filters.
        8 - HVAC radiation around 43 MHz
John reported that the servo room shielding and filtering has been tightened up, and more work needs to be done. A number of unfiltered wires were found to be going through the room walls. A generic filtered connector is needed for getting signals into and out of this room. A strong 300.00 MHz signal from the telephone system was seen.
        9 - DGS laser system data acquisition module
No report on this area.
        10 - 370 MHz wideband noise toward Arbovale (Jeff, Wes)
Rick showed an azimuth sweep of the deer stand antenna that indicates two or three sources of broadband noise coming from the general direction of Arbovale. It is strongest over a few MHz around 370 MHz. Wes and Jeff will look into it with the truck.
        11 - Blue warning lamps (John)
John reported on discussions with Pete Chestnut on the need for the blue warning lights on the GBT and other lamps that may need to be replaced because they use motors with brushes. Agreement on required changes are nearly in hand.
        12 - Main GBT drives (Rick)
Rick measured one GBT azimuth slew with the deer stand between about 350 and 370 MHz. The strongest RFI appeared around times of maximum acceleration and deceleration of the antenna, but it appears that there is broadband radiation during both tracking and slewing. A better measurement setup is required because the deer stand is partially shielded from the base of the GBT by a grove of pine trees. John emphasized the need to determine whether any major modifications to the drives are needed fairly soon so that this work can be accomplished during the anticipated shutdown this summer for track grout fixes. John, Randy, and Rick will assemble a group to attack this problem.
        13 - Anechoic chamber turntable (Jeff, Carla)
Carla reported that the Scientific Atlanta rotator works. She and Jeff are looking into rotary joints for RF and power lines and are in the process of drawing up requirements. Steve White has been looking for rotary joints for the outdoor antenna range so he will be consulted. Carla is working with Mike on the need to excavate the anechoic chamber's concrete floor to make the RFI measurement turntable low enough to avoid interference with the antenna pattern measurements.
        14 - Call-out procedures and assignments (Jeff, John)
Wes will be the primary contact for GBT interference reports from observers and operators. Jeff and Carla will shadow Wes for awhile to pick up on the techniques for responding to RFI reports. Jeff will give a call-out procedure to Pete and the contact list on the web will be modified to include RFI call-outs.
        15 - Site-wide interference suppression policy status (John)
No progress.
        16 - RFI considerations for new Education Center (John, Frank)
John expects a meeting to take place soon.
        17 - Equipment inventory (Jeff)
Jeff reported that an HP-8340B synthesizer/sweeper is now available at the anechoic chamber. A network analyzer head for this unit may also be available.
        18 - Deer stand RFI survey (Rick)
Nothing to report
        17 - GBT RFI 'harmful levels' document and measurement   procedures
Not discussed
        18 - Power line noise monitoring
Not discussed
        19 - Measurements of PC/Workstation in a shielded box
Not discussed
        20 - Interference at 1612.17 MHz
Not discussed
        21 - New and other action items
Atmospheric phase interferometer (Carla): Some progress made on suppression.

CenterComm unit (Carla): A shielded box is in the works.

A main AC power switch is need for the auditorium projection rack. (Carla)

Storage area: Wes pointed out the need for a storage room for IPG equipment. He will look into using one or more areas in the interferometer control room. Jeff noted to need to sort out and store some equipment in the barn near the school.

    Initiatives to assemble requirements and specifications in the following areas:
Not discussed
 Next Meeting:  4 - 16 - 02, 0800